University of New Mexico Department of Computer Science

Early Course Evaluation

CS 591 Randomized Algorithms Fall Semester, 2004

Please help me make the course better by answering the following short questions.

How clear are the lectures?  1 2 3 4 5 6
Not clear  Very clear
How is the pace of the class?  1 2 3 4 5 6
Too slow  Too fast
How is the difficulty level of the material?  1 2 3 4 5 6
Too easy  Too hard
How is the difficulty level of the homeworks?  1 2 3 4 5 6
Too easy  Too hard
Do the homeworks enhance your understand-  1 2 3 4 5 6
ing of the material? Not at all Very much

What is the single best thing that can be done to improve the quality of the class? What is the most interesting thing you’ve learned in the class up to this point? What is the most confusing thing you’ve learned in the class up to this point? Other comments, suggestions, ideas, complaints, compliments?